OpenStudy (wonderwomen):

what is the sum in its simplest form? 11/14 + 37/42=?

OpenStudy (anonymous):

Both 14 and 42 have a factor in common: 7. What happens if you multiply 11/14.3/3?

OpenStudy (anonymous):

\[\frac{ 11 }{ 14 }.\frac{ 3 }{ 3 }=?\]

OpenStudy (wonderwomen):

you get 33/42

OpenStudy (anonymous):

Yes... now you can add \[\frac{ 33 }{ 42 }+\frac{ 37 }{ 42 }=?\]

OpenStudy (wonderwomen):

70/42 but you have to simplfy and you get 1 and 2/3

OpenStudy (anonymous):

70/42 is correct, and yes you have to simplify. Divided by 7 it becomes 10/6. Divided by 2 it becomes..?

OpenStudy (wonderwomen):

5/3 but you have to get it into a proper fraction and then it goes to 1 and 2/3

OpenStudy (anonymous):

1+2/3 is correct, but it's better to keep it as one fraction

OpenStudy (anonymous):

Well you know what your teacher wants. As a teacher I want my pupils to keep it as 5/3.